# Climate Science and Control Engineering: Insights, Parallels, and Connections

### v2

@@ Abstract
Climate science is the multidisciplinary field that studies the Earth’s climate and its evolution—past, present, and future. This simultaneously includes studying the natural and human-caused factors that meaningfully influence and often adversely impact it. At the very core of climate science are indispensable climate models that predict future climate scenarios, inform policy decisions, and dictate how a country’s economy should change in light of the changing climate. Climate models capture a wide range of interacting dynamic processes via extremely complex ordinary and partial differential equations—processes that span varying spatiotemporal scales. To model these large-scale complex processes, state-of-the-art climate science leverages supercomputers, advanced simulations, statistical methods, and machine learning techniques to predict future climate.
@@
Index Terms:
Climate modeling, control engineering, dynamic systems, climate-driven decision making, climate mitigation strategies.
@@ I Introduction and Tutorial Objectives
Despite significant advances in computational power, statistical and stochastic methods, and machine learning integration, control engineering principles are rarely presented as an alternative to study climate-related problems. Climate systems are at their core dynamic in nature, governed by underlying physical laws and exhibiting complex behaviors that evolve over time. However, the application of control engineering principles (which are fundamentally suited to analyzing and managing such dynamic systems) remains relatively underexplored in the field of climate science. The objectives of this tutorial are three-fold: (i) introduce control engineering researchers to climate science and modeling; (ii) produce a fresh outlook on climate problems; and (iii) build parallels between various known and well-studied climate science problems and their control theoretic counterparts.
Figure 1: Climate as a dynamic system. In this tutorial, we explain how the Earth’s climate can be viewed as a complex dynamical system with inputs, outputs, internal states, and external disturbances. We describe how the key components of the climate system—atmosphere, hydrosphere, cryosphere, biosphere, and land surface—interact and evolve over time. We cover the roles of natural and anthropogenic inputs (e.g., solar radiation, green infrastructure, and natural variability) and disturbances (e.g., emissions and deforestation), and how they influence the system’s trajectory. We delineate the different types and levels of climate models and highlight how those models map the climate inputs and disturbances to evolving states and observable outputs such as global average temperature, in the form of scenario-based simulations and projections. Throughout the tutorial, we highlight and examine these dynamic processes through a control-theoretic lens and discuss how control tools can provide structure for understanding, simulating, and potentially influencing climate dynamics.
@@
Some Efforts in Controls for Climate. While some efforts have been made to apply optimization and control techniques to climate-related challenges, they have largely focused on specific goals or methodologies distinct from the approach presented here. For example, a group of researchers has examined weather and climate systems as dynamic systems influenced by human activity, with a particular emphasis on aerosol injections. In the studies [6, 7], the authors formulated optimal control problems to explore deliberate weather and climate modification. The authors modeled the atmosphere as a closed-loop dynamical system, where state variables included atmospheric parameters such as temperature and pressure, while control variables represented human interventions like aerosol injections. Their objective was to modify natural atmospheric processes to achieve desired weather outcomes, framed as a dynamic optimization problem. In their most recent work [8], they applied optimal control theory to develop strategies for stratospheric aerosol injections, aiming to stabilize global climate temperatures by adjusting aerosol emission rates.
@@
Tutorial Objectives and Broader Implications. This tutorial aims to build a foundational framework that connects the principles of control engineering with the complexities of climate science by:
  • Control-Theoretic Insights and Applications: We survey and demonstrate the relevance, potential, and applicability of control-theoretic and network-based techniques that can be applied to the control-oriented representations of climate dynamics. These techniques are further explored for their implications and outcomes in developing physics-informed frameworks for climate learning, control, and estimation. Leveraging such frameworks supports decision-making processes and drives mitigation efforts. Additionally, we identify potential research directions and practical applications that can benefit from this interdisciplinary approach. (Sections V, VI, and VII)
  • Drawing Parallels Between Climate Science Challenges and Control Engineering Concepts: We highlight conceptual and structural similarities between problems in climate science and control engineering. These include, for instance, the role of observability in inferring unmeasured climate states from sparse data. We also discuss the resemblance between climate model calibration and system identification, the parallels between uncertainty quantification in climate projections and robust control under model uncertainty, and the interpretation of mitigation policies as constrained optimal control problems. We note that although we make these observations in this manuscript, the aforementioned problems have not been studied through the lens of control engineering. By outlining these connections, we aim to foster a shared vocabulary and problem structure that can bridge the two fields and promote new interdisciplinary strategies. (Sections V, VI, and VII)
It is important to emphasize that the purpose of this tutorial is not to teach climate scientists control theory, nor to suggest that existing climate science lacks methodological depth. Rather, this tutorial is intended for the control engineering community, aiming to introduce climate systems as a rich, complex, and practically relevant domain where control-theoretic tools can be meaningfully applied. Climate models provide natural testbeds for developing and evaluating methods in large-scale dynamical systems, estimation under uncertainty, and robust decision-making—all areas where control theory has much to contribute. We simply describe what control systems theory can potentially deliver for climate research.
The structure of the tutorial is designed to support this goal and the aforementioned objectives: it begins with the basics and foundations of climate science and dynamics (Section II). Section III explains the evolution and classification of climate models. Section IV discusses the distinction between weather and climate and how these concepts can be interpreted within the control-oriented framework introduced in this tutorial. Section V highlights the dynamic nature of climate models from a control-theoretic perspective and poses questions that can be tackled using control methods. This is followed by Section VII, which examines climate dynamics through the lens of control engineering, offering interpretations and insights. In particular, this section draws explicit parallels between challenges in climate science and their counterparts in control theory. Section VIII concludes the tutorial with its summary and potential next steps and recommendations for the interested researchers.
@@
Figure 3: Climate system’s main five components.
@@ II Birds’ Eye View of Climate Models: From Simple Physics to the Discipline of Earth Science
III Climate Models: Evolution, Scales, and Applications
III-A Evolution and Classification of Climate Models
@@
Figure 6: Figures generated by the IPCC AR6 Atlas [122], showing projected changes in mean annual temperature (°C) relative to the 1850–1900 baseline under different emission scenarios leading to average global warming of $1.5^{\circ}$, $2^{\circ}$, $3^{\circ}$, and $4^{\circ}$. Solid areas indicate high agreement among CMIP6 models (34 models for the first three scenarios, 20 for the last), while hatched areas indicate low agreement.
@@
An Example on Predicting and Computing Climate Projections. Here, we illustrate how climate projections—that have become part of the climate discourse—are computed via dynamic climate models. For example, Fig. 6 presents visualizations generated by the IPCC’s Interactive Atlas, displaying projected changes in mean annual temperature relative to the 1850–1900 baseline [^1] The 1850–1900 baseline is used by the IPCC as a proxy for pre-industrial conditions, as it represents the average temperature before significant industrialization and the associated increase in GHGs emissions. It is computed by averaging observed or reanalyzed surface temperatures over that period. While direct instrumental coverage was sparse, especially in the Southern Hemisphere, recent studies [123, 124] show that the baseline is accurate within about ±$\pm 0.1^{\circ}$, making it suitable for long-term comparisons and global warming targets. under four global warming levels: $1.5^{\circ}$, $2^{\circ}$, $3^{\circ}$, and $4^{\circ}$. These specific levels represent policy-relevant thresholds and scientific milestones in the context of climate change mitigation and adaptation. The $1.5^{\circ}$ and $2^{\circ}$ targets are directly tied to the Paris Agreement, [125]—an international treaty adopted under the United Nations Framework Convention on Climate Change (UNFCCC) that aims to limit global warming to well below 2C, with efforts to pursue a 1.5C target above pre-industrial levels. Each additional degree of warming brings increasingly severe consequences. That being said, the $3^{\circ}$ and $4^{\circ}$ levels serve as illustrative benchmarks for high-emission scenarios. These levels are also widely used in impact assessments across various domains, such as ocean science, water availability, energy systems, agriculture, and public health, where researchers estimate physical and societal impacts under different degrees of warming. This makes them a common language for linking model projections to real-world consequences.
@@
These projections are derived from a multi-model ensemble of CMIP6 simulations, using 34 dynamic models for the first three scenarios and 20 for the fourth. The reduced number of models at the $4^{\circ}$ warming level is primarily due to the fact that not all modeling centers have submitted simulations for high-emission or extended scenarios, and some models were excluded due to missing output fields or inconsistencies that affect comparability. From a systems and dynamics perspective, each of these models solves a set of coupled nonlinear ODEs and PDEs that govern the evolution of climate variables such as temperature, wind, ocean currents, and moisture. These PDEs are discretized over spatial grids and time steps, often with horizontal resolutions ranging from tens to hundreds of kilometers, which means that generating a single simulation involves high computational cost and long runtimes. For example, simulating a century-long climate trajectory at high spatial and temporal resolution requires significant computational infrastructure, limiting how many runs some centers can produce, especially for extreme scenarios. The plots shown are constructed by spatially averaging the simulated surface temperature fields over time for each model grid, followed by computing the ensemble mean across all models included for a given warming level. Agreement and confidence are visually encoded: solid-shaded regions indicate strong agreement among the models (i.e., consistent sign of change in at least 80% of models), while hatched regions indicate areas of low confidence where models disagree on the direction of change or the signal-to-noise ratio is weak.
@@
The Interactive Atlas offers additional features for other climate variables such as surface wind and precipitation, and enables users to generate regional time series, assess seasonal variability, and explore different emissions scenarios through an interactive interface. Tools such as AR6 Atlas and the CMIP6 simulations are a rich source that can be used as testbed for generating data that can be used to solve climate-relevant control engineering problems as emphasized in Section VII.
@@ III-B How are Climate Models Useful?
The concepts of weather and climate are sometimes confused and used interchangeably. Weather refers to the short-term atmospheric conditions in a specific place at a specific time. It encompasses daily variations in temperature, humidity, precipitation, wind, and visibility, along with severe conditions including droughts, tornadoes, and hurricanes. These conditions can change within minutes or hours. In contrast, climate represents the long-term average of weather patterns over an extended period—months, years, or even centuries. As NOAA puts it: Climate is what you expect, weather is what you get.
From a control systems perspective, this distinction is analogous to analyzing a system’s short-term response (transients) versus its long-term behavior (average or steady-state response). While weather prediction involves tracking the immediate state trajectory of a highly nonlinear, chaotic system, climate modeling focuses on its statistical or average behavior over time making concepts related to infinite-horizon or asymptotic stability, reachability, and observability relevant for climate dynamics. This shift in focus from state evolution to long-term system characteristics is key for developing control-oriented strategies for climate adaptation and mitigation.
@@ IV-A Entities Providing Weather Forecasts and Climate Projections
Why is this information about these meteorological agencies and their forecasting models important for the controls community? Many of these agencies provide open access to full or partial datasets of historical and real-time weather data for academic research and scientific applications. Additionally, analyzing the differences in their model structures, time-frames, and methodologies and linking them to control engineering dynamics and applications discussed in this tutorial creates opportunities for further research and practical implementations.

### v3

@@ Abstract
Climate science is the multidisciplinary field that studies the Earth’s climate and its evolution—past, present, and future. This simultaneously includes studying the natural and human-caused factors that meaningfully influence and often adversely impact it. At the very core of climate science are indispensable climate models that predict future climate scenarios, inform policy decisions, and dictate how a country’s economy should change in light of the changing climate. Climate models capture a wide range of interacting dynamic processes via extremely complex ordinary and partial differential equations—processes that span varying spatiotemporal scales. To model these large‑scale complex processes, state-of-the-art climate science leverages supercomputers, advanced simulations, statistical methods, and machine learning techniques to predict future climate.
@@
SAI Stratospheric Aeorosal Injection
SRM Solar Radiation Management
@@ I Introduction and Tutorial Objectives
Despite significant advances in computational power, statistical and stochastic methods, and machine learning integration, control engineering principles are rarely presented as an alternative to study climate-related problems. Climate systems are at their core dynamic in nature, governed by underlying physical laws and exhibiting complex behaviors that evolve over time (Fig. 1). However, the application of control engineering principles (which are fundamentally suited to analyzing and managing such dynamic systems) remains relatively underexplored in the field of climate science. The objectives of this tutorial are three-fold: (i) introduce control engineering researchers to climate science and modeling; (ii) produce a fresh outlook on climate problems; and (iii) build parallels between various known and well-studied climate science problems and their control theoretic counterparts.
Figure 1: Climate as a dynamic system. In this tutorial, we explain how the Earth’s climate can be viewed as a complex dynamical system with inputs, outputs, internal states, and external disturbances. We describe how the key components of the climate system—atmosphere, hydrosphere, cryosphere, biosphere, and land surface—interact and evolve over time. We cover the roles of natural and anthropogenic inputs (e.g., solar radiation, green infrastructure, and natural variability) and disturbances (e.g., emissions and deforestation), and how they influence the system’s trajectory. We delineate the different types and levels of climate models and highlight how those models map the climate inputs and disturbances to evolving states and observable outputs such as global average temperature, in the form of scenario-based simulations and projections. Throughout the tutorial, we highlight and examine these dynamic processes through a control-theoretic lens and discuss how control tools can provide structure for understanding, simulating, and potentially influencing climate dynamics. We note here, as a much-needed disclaimer, that jointly modeling all of the dynamic system components and inputs/outputs in this figure is virtually impossible as such models would not be practical computationally. Climate models wisely stick to subsets of these interactions rather than the whole kitchen sink, and we, too, respect that wisdom (mostly). This tutorial offers only some perspectives that are neither complete nor extraordinary. It just makes climate science more accessible to control system researchers and hopes for more cross-disciplinary learning and developments.
@@
Some Efforts in Controls for Climate. While some efforts have been made to apply optimization and control techniques to climate-related challenges, they have largely focused on specific goals or methodologies distinct from the approach presented here. For example, a group of researchers has examined weather and climate systems as dynamic systems influenced by human activity, with a particular emphasis on aerosol injections. In the studies [6, 7], the authors formulated optimal control problems to explore deliberate weather and climate modification. The authors modeled the atmosphere as a closed-loop dynamical system, where state variables included atmospheric parameters such as temperature and pressure, while control variables represented human interventions like aerosol injections. Their objective was to modify natural atmospheric processes to achieve desired weather outcomes, framed as a dynamic optimization problem. In their most recent work [8], they applied optimal control theory to develop strategies for stratospheric aerosol injections, aiming to stabilize global climate temperatures by adjusting aerosol emission rates. Another recent paper [9] formally defines control inputs within climate models and demonstrates how scenario generation for aerosol injection can be treated as a control-guided dynamic process. In parallel, the work of [10] frames geoengineering as a multivariable feedback control problem, emphasizing the design of climate intervention strategies that explicitly account for system uncertainty and feedback latency via a simple control methodologies. This design-driven view builds on physical and spatial insights from [11] and [12], who explore the spatial heterogeneity and hemispheric responses to solar radiation management.
@@
Tutorial Objectives and Broader Implications. This tutorial aims to build a foundational framework that connects the principles of control engineering with the complexities of climate science.
Tutorial Goal: Bridge Control and Climate Science
$\bullet$ Establish a framework linking control and climate research.
$\bullet$ Produce brief intro to various dynamic climate models.
$\bullet$ Summarize control efforts in climate literature.
$\bullet$ Build parallels between climate and control problems.
$\bullet$ Formulate novel high-level climate control problems.
(Sec. IIIII) Introduce Climate Components
$\bullet$ Discuss history of climate models and explain key physical processes & dynamic nature of climate systems.
$\bullet$ Present the applicability and goal of climate models.
$\bullet$ Review model hierarchies, spatiotemporal scales, and limitations.
(Sec. IVVI) Control Modeling for Climate Problems
$\bullet$ Recast climate models as MIMO state-space systems.
$\bullet$ Describe rudimentary and more complex climate models.
$\bullet$ Identify states, inputs, uncertainty, and measurable outputs.
$\bullet$ Facilitate estimation, control, uncertainty quantification.
(Sec. IVVI) Control-Theoretic Insights & Applications
$\bullet$ Demonstrate role of control for climate regulation.
$\bullet$ Describe observability and controllability-related problems in climate dynamics.
$\bullet$ Explore control-based, geoengineering design frameworks for mitigation and adaptation strategies.
(Sec. VIII) Parallels Between Climate & Control Sciences
$\bullet$ Observability analysis and state estimation $\leftrightarrow$ state inference from sparse data and missing data problem.
$\bullet$ System identification $\leftrightarrow$ climate model calibration.
$\bullet$ Optimal control $\leftrightarrow$ mitigation and adaptation policies.
(Sec. VIII) Open Problems in Climate Science:
$\bullet$ Inferring hidden states from sparse/noisy observations.
$\bullet$ Quantifying uncertainty in nonlinear, large-scale models.
$\bullet$ Designing robust and optimal mitigation/control policies.
$\bullet$ Integrating physics-based and data-driven learning frameworks.
(Sec. IX) Tutorial Summary and Concrete Problems:
$\bullet$ Climate models as testbeds for advancing control theory.
$\bullet$ Reachability & observability of large-scale climate models.
$\bullet$ Optimized sensor placement and orbital designs.
$\bullet$ Scalable model order reduction and system identification.
Figure 3: Tutorial summary and paper organization.
  • Control-Theoretic Insights and Applications: We survey and demonstrate the relevance, potential, and applicability of control-theoretic and network-based techniques that can be applied to the control-oriented representations of climate dynamics. These techniques are further explored for their implications and outcomes in developing physics-informed frameworks for climate learning, control, and estimation. Leveraging such frameworks supports decision-making processes and drives mitigation efforts. Additionally, we identify potential research directions and practical applications that can benefit from this interdisciplinary approach. Some control engineering efforts have taken place under the umbrella of climate geoengineering. Such efforts are briefly summarized too. (Sections V, VI, VII, and VIII)
  • Drawing Parallels Between Climate Science Challenges and Control Engineering Concepts: We highlight conceptual and structural similarities between problems in climate science and control engineering. These include, for instance, the role of observability in inferring unmeasured climate states from sparse data. We also discuss the resemblance between climate model calibration and system identification, the parallels between uncertainty quantification in climate projections and robust control under model uncertainty, and the interpretation of mitigation policies as constrained optimal control problems. We note that although we make these observations in this manuscript, the aforementioned problems have not been studied through the lens of control engineering. By outlining these connections, we aim to foster a shared vocabulary and problem structure that can bridge the two fields and promote new interdisciplinary strategies. (Sections V, VI, and VIII)
It is important to emphasize that the purpose of this tutorial is not to teach climate scientists control theory, nor to suggest that existing climate science lacks methodological depth. Rather, this tutorial is intended for the control engineering community, aiming to introduce climate systems as a rich, complex, and practically relevant domain where control-theoretic tools can be meaningfully applied. Climate models provide natural testbeds for developing and evaluating methods in large-scale dynamical systems, estimation under uncertainty, and robust decision-making—all areas where control theory has much to contribute. We simply describe what control systems theory can potentially deliver for climate research. The structure of the tutorial is designed to support this goal and the aforementioned objectives. The tutorial organization is provided in Fig. 3.
@@ II Birds’ Eye View of Climate Models: From Simple Physics to the Discipline of Earth Science
AtmosphereGases & WindsAtmosphereGases & WindsHydrosphereCurrents & SalinityCryosphereIce & SnowLandSoil & VegetationBiosphereLife & BiomassHeat FluxRadiation TempEvapotransp.Precipitation TempNutrients HeatWater CarbonMelting FreshwaterCO2 releaseEvapotranspirationHeat Flux
Figure 4: Climate system’s main five components and selected interactions between them (illustrative, not exhaustive).
@@
III Climate Models: Evolution, Scales, and Applications
III-A Evolution and Classification of Climate Models
@@
Figure 7: Figures generated by the IPCC AR6 Atlas [126], showing projected changes in mean annual temperature (°C) relative to the 1850–1900 baseline under different emission scenarios leading to average global warming of $1.5^{\circ}$C, $2^{\circ}$C, $3^{\circ}$C, and $4^{\circ}$C. Solid areas indicate high agreement among CMIP6 models (34 models for the first three scenarios, 20 for the last), while hatched areas indicate low agreement.
@@
An Example on Predicting and Computing Climate Projections. Here, we illustrate how climate projections—that have become part of the climate discourse—are computed via dynamic climate models. For example, Fig. 7 presents visualizations generated by the IPCC’s Interactive Atlas, displaying projected changes in mean annual temperature relative to the 1850–1900 baseline [^1] The 1850–1900 baseline is used by the IPCC as a proxy for pre-industrial conditions, as it represents the average temperature before significant industrialization and the associated increase in GHGs emissions. It is computed by averaging observed or reanalyzed surface temperatures over that period. While direct instrumental coverage was sparse, especially in the Southern Hemisphere, recent studies [127, 128] show that the baseline is accurate within about $\pm 0.1^{\circ}$C, making it suitable for long-term comparisons and global warming targets. under four global warming levels: $1.5^{\circ}$C, $2^{\circ}$C, $3^{\circ}$C, and $4^{\circ}$C. These specific levels represent policy-relevant thresholds and scientific milestones in the context of climate change mitigation and adaptation. The $1.5^{\circ}$C and $2^{\circ}$C targets are directly tied to the Paris Agreement, [129]—an international treaty adopted under the United Nations Framework Convention on Climate Change (UNFCCC) that aims to limit global warming to well below 2C, with efforts to pursue a 1.5C target above pre-industrial levels. Each additional degree of warming brings increasingly severe consequences. That being said, the $3^{\circ}$C and $4^{\circ}$C levels serve as illustrative benchmarks for high-emission scenarios. These levels are also widely used in impact assessments across various domains, such as ocean science, water availability, energy systems, agriculture, and public health, where researchers estimate physical and societal impacts under different degrees of warming. This makes them a common language for linking model projections to real-world consequences.
@@
These projections are derived from a multi-model ensemble of CMIP6 simulations, using 34 dynamic models for the first three scenarios and 20 for the fourth. The reduced number of models at the $4^{\circ}$C warming level is primarily due to the fact that not all modeling centers have submitted simulations for high-emission or extended scenarios, and some models were excluded due to missing output fields or inconsistencies that affect comparability. From a systems and dynamics perspective, each of these models solves a set of coupled nonlinear ODEs and PDEs that govern the evolution of climate variables such as temperature, wind, ocean currents, and moisture. These PDEs are discretized over spatial grids and time steps, often with horizontal resolutions ranging from tens to hundreds of kilometers, which means that generating a single simulation involves high computational cost and long runtimes. For example, simulating a century-long climate trajectory at high spatial and temporal resolution requires significant computational infrastructure, limiting how many runs some centers can produce, especially for extreme scenarios. The plots shown are constructed by spatially averaging the simulated surface temperature fields over time for each model grid, followed by computing the ensemble mean across all models included for a given warming level. Agreement and confidence are visually encoded: solid-shaded regions indicate strong agreement among the models (i.e., consistent sign of change in at least 80% of models), while hatched regions indicate areas of low confidence where models disagree on the direction of change or the signal-to-noise ratio is weak.
@@
The Interactive Atlas offers additional features for other climate variables such as surface wind and precipitation, and enables users to generate regional time series, assess seasonal variability, and explore different emissions scenarios through an interactive interface. Tools such as AR6 Atlas and the CMIP6 simulations are a rich source that can be used as testbed for generating data that can be used to solve climate-relevant control engineering problems as emphasized in Section VIII.
@@ III-B How are Climate Models Useful?
The concepts of weather and climate are sometimes confused and used interchangeably. Weather refers to the short-term atmospheric conditions in a specific place at a specific time. It encompasses daily variations in temperature, humidity, precipitation, wind, and visibility, along with severe conditions including droughts, tornadoes, and hurricanes. These conditions can change within minutes or hours. In contrast, the atmosphere component in climate represents the long-term average of weather patterns over an extended period—months, years, or even centuries. As NOAA puts it: Climate is what you expect, weather is what you get.
From a control systems perspective, this distinction is analogous to analyzing a system’s short-term response (transients) versus its long-term behavior (average or steady-state response). While weather prediction involves tracking the immediate state trajectory of a highly nonlinear, chaotic system, climate modeling focuses on statistical or average behavior over time, making concepts like infinite-horizon or asymptotic stability, reachability, and observability relevant. This shift in focus from state evolution to long-term system characteristics is key for developing control-oriented strategies for climate adaptation and mitigation.
@@ IV-A Entities Providing Weather Forecasts and Climate Projections
Why does information about meteorological agencies and their forecasting models matter to the controls community? Many of these agencies provide open access to full or partial datasets of historical and real-time weather data for academic research and scientific applications. Additionally, analyzing the differences in their model structures, time-frames, and methodologies and linking them to control engineering dynamics and applications discussed in this tutorial creates opportunities for further research and practical implementations.